OpenStudy (anonymous):

Simplify square root of 80 end root minus 3 square root of 63 end root square root of 125 minus square root of 175

OpenStudy (anonymous):

\[\sqrt{80}-(3\sqrt{63})(\sqrt{125})-\sqrt{175}\] like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first simplify each radical.. for example